what is 3/4 minus 5/7 in lowesr terms
Found 2 solutions by user_dude2008, rfer:
Answer by user_dude2008(1862)

3/4-5/7 = 21/28-20/28 = (21-20)/28 = 1/28

3/4-5/7 = 1/28

Answer by rfer(16322)

21/28-20/28=1/28